OpenStudy (shortstacc):

Suppose an investment of $3,800 doubles in value every decade. The function gives the value of the investment after x decades. How much is the investment worth after 2 decades?

OpenStudy (tkhunny):

Doubles every decade. Two decades. Please provide some idea.

OpenStudy (danjs):

It will be an exponential function y = a*b^x every decade increase in x, the value of y doubles or is multiplied by 2. and starts at 3800 (x , y) = (0 , 3800) y = a*b^x 3800 = a*b^0 = a*1 a=3800 y = 3800*b^x Each decade for x increases by 1, the 3800 is multiplied by 2. y =3800 * 2^x x is the number of decades and y is the total investment